Solve the following equations in the interval [0,2].
Note: Give the answer as a multiple of pi. Do not use decimal numbers. The answer should be a fraction or an integer. Note that pi is already included in the answer so you just have to enter the appropriate multiple. E.g. if the answer is pi/2 you should enter 1/2. If there is more than one answer enter them separated by commas.
sin(t) = -sqrt3/2; t = ? pi
sin(t) = 1/2; t = ? pi

Answer by Alan3354(21605)

sin(t) = -sqrt3/2; t = ? pi
t = 4pi/3
t = 5pi/3
------------------
sin(t) = 1/2; t = ? pi
t = pi/6
t = 5pi/6